V(4)= 4/3π R^3
Help I just don't understand how to do this.

Answer by stanbon(75887)

V(4)= 4/3π R^3
Help I just don't understand how to do this.
---------------------------
Formula:
V(r) = (4/3)(pi)r^3
---------------------------
v(4) = (4/3)(pi)4^3
V(4) = (4/3)(pi)64
V(4) = 268.08
===================
Cheers,
Stan H.